Shift summary owl 7/29/99

Mary_Manning (
Thu, 29 Jul 1999 07:16:21 -0500

Shift Summary for owl shift, 7/29/99

On Shift: Val Prasad, Lisa Manning (Guest Appearance: E Blucher)

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E832 data (special runs not included): 4.9
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time: drt_fresh_starts, recovering from ktev4
failure, kfrends


Runs ended this shift: 14051

Tapes brought to FCC this shift: QK*188-190, QKB 064


Accesses during shift:none


Overview of shift: Smooth 832 data-taking until two dead L3 CPU
processes in ktev1 and ktev2 necessitate a drt_fresh_start. Ktev4
crashes while the run is being stopped, so new tapes have to be loaded.

Also, since ktev4 failed, we must run in BEAM_832_NOKTEV4
configuration. Smooth 832 data-taking for the remainder of the shift.


Description of Shift:

run spill time comment

14051 00:00 Take over from evening shift

139,150 ~00:50 Ktevana errors. Solved by kfrend.

182 01:35 Ktev2 process 16 and Ktev1 process 4 are dead.
Livetime falls drastically.

01:40 Stop run 14051 to do a drt_fresh_start.
While run is being stopped, ktev4 stops
responding and run does not end sucessfully.
Histograms are saved,though.
Tapes QK*188-190 and QKB 064 are removed.

01:50 Found that ktev4 does not have power, and
power cycling does not help. The power switch
is left in off position.

We modified various files(see logbook) to
run in BEAM_832_NOKTEV4 configuration.

DDD Veto and EOS Spill signals to KTEV4 pulled
out of 4616 in Rack RR168 (Look for the yellow
labels !)

After several unsucessful inits and
drt_fresh_starts, run sucessfully initialized.

14057 3:55 Start 832_beam_NOKTEV4 run

5:30 Notified help desk of problem with ktev4.


Follow up Notes on Previous Shift: --


Items to Follow up on Next Shift:
Do not use ktev4 until it has been repaired. Continue with
BEAM_832_NOKTEV4 configuration.
Follow up on request to get KTEV4 fixed.